OpenStudy (anonymous):

if m^15 = 48/y and m^13 = 2/6y and m>0 what is the value of m? show work and there are no choices

OpenStudy (campbell_st):

well look at it this way \[m^{15} =m^2 \times m^{13}\] so doing some substitution \[m^{15} = m^2 \times \frac{2}{6y}\] which becomes \[\frac{48}{y} = m^2 \times \frac{2}{6y}\] divide both sides by 2/6y and you get \[\frac{48}{y} \times \frac{6y}{2} = m^2\] this should allow you to find m

OpenStudy (anonymous):

so then i cross multiply to find y then plug y in to find m?

OpenStudy (campbell_st):

well looking at the 2 fractions on the left, cancel common factors and you  evaluate then find the value of m